Skip to end of metadata
Go to start of metadata

We met with the tower company again today and appear to have a path forward to lift our climbing stoppage.  The company will fix the saddles, when obvious problems are noted.  They will retension towers with us monitoring with our tensiometers (these should be in FedEx to us now, to arrive Wed.).  Clearly this will take some time, but is the solution we wanted.

Today's tasks:

  • Prepped tse05 (the last 60m tower) for the INEGI crew to install tomorrow.  They installed 2 of the 3 solar panels on this tower today to get ready.
  • Connected AC power to rne02 and rne03, which are now reporting (though rne02's 20m RMY is not reporting good data).
  • More ops center network work.  We did have a brief outage that took a while to recover from since I let a key go up the mountain.
  • Prepped the remaining (all?) user sensor booms

DTU and INEGI crews are now at the site, ARL having left over the weekend.  Cornell is still here as well.

 

 

 

  • No labels